# abEunwoo Heo    abByeongchan Choi    abJae-Hun Jung Corresponding author

### v1

@@ Abstract
Topological data analysis over graph has been actively studied to understand the underlying topological structure of data. However, limited research has been conducted on how different distance definitions impact persistent homology and the corresponding topological inference. To address this, we introduce the concept of path-representable distance in a general form and prove the main theorem for the case of cost-dominated distances. We found that a particular injection exists among the $1$-dimensional persistence barcodes of these distances with a certain condition. We prove that such an injection relation exists for 00- and $1$-dimensional homology. For higher dimensions, we provide the counterexamples that show such a relation does not exist.
@@ 1 Introduction
Topological data analysis (TDA), developed recently, has proven its usefulness in various applications [1, 2, 3, 4]. Persistent homology (PH) is one of the primary tools in TDA, using homological characteristics of the data at various scales [2]. The PH method involves constructing the complex in different scales sequentially, capturing how homology changes as the complex is being built and refined. This procedure is known as filtration. This hierarchical changes of homology enables the PH method to provide the topological inference of the data.
@@
Once the appropriate transform is identified, applying the PH method requires defining a proper metric for the transformed geometric object, determining the distance between points in the point cloud or between nodes in the graph representation. There is no single way to define this metric or distance as there can be multiple definitions. Definitions of such distances have been empirically established in different fields of study, yet there is a lack of research on how these varying definitions of distance impact PH calculations.
@@
On the other hand, the shortest path as the path connecting two vertices with the least number of edges can be also considered. In a similar sense, there are data in which the relevancy between two nodes is more explained in the path passing through the minimum number of edges than in that with minimum sum of weights. In such a case, it is natural to define the distance between two vertices using such paths. The first approach was conducted in previous research [15]. We canonically reformulated the definition of distance $d_{\text{edge}}(v,w)$ between two points on the graph to be the minimum sum of weights between the paths connecting $v$ and $w$ with the minimum number of edges.
@@
In this research we found an interesting relation between $d_{\text{weight}}$ and $d_{\text{edge}}$. That is, for $d_{\text{weight}}\leq d_{\text{edge}}$, we have $\textbf{bcd}_{1}(d_{\text{weight}})\subseteq\textbf{bcd}_{1}(d_{\text{edge}})$, where $\textbf{bcd}_{1}$ is the $1$-dimensional barcode. This property is quite unusual and is generally not expected to hold, as depicted in Figure 1. To explore this, we introduce a more general concept of distance, termed path-representable distance.
@@
Furthermore in this paper, we prove that such an injection relation holds among any path-representable distances for $1$-dimensional homology. This paper is composed of the following sections. In Section 2, explain distinct distance definitions over the given graph, particularly the weight- and edge-minimal distances and the possible relations among them. In Section 3, we introduce the concept of path-representable distance, a generalization of $d_{\text{weight}}$ and $d_{\text{edge}}$ and we introduce the main theory related to path-representable distances. In Section 4, we rigorously prove the main propositions and theorem. In Section 5, we make an experiment to explain how the main theorem works on the higher dimensions. Finally, in Section 6, we provide a brief concluding remark.
@@ 2.1 Different distance definitions
$d_{\text{edge}}(v,w)=\min_{p}\left\{\sum_{e\in E(p)}W_{E}(e)\mathrel{\bigg{|}}p\text{ is a path between }v\text{ and }w\text{ such that }|E(p)|=m_{vw}\right\},$
@@ 2.2 Persistent homology
for any $k$-simplex $\sigma\in\mathbb{X}$ whenever $p\geq 1$, otherwise $f(\sigma)=0$. For a sequence of real numbers $0=a_{0}\leq a_{1}\leq\ldots\leq a_{m}$, if we denote $\mathbb{X}_{a_{l}}\coloneqq f^{-1}((-\infty,a_{l}])$, then we have a filtration
@@
Note that $\mathbb{X}_{a_{0}}=V$ and $\mathbb{X}_{a_{m}}=\mathcal{P}(V)$. This filtration is called the Vietoris-Rips (Rips) filtration. The filtration induced by $f$ allows us to study topological changes in the simplicial complex $\mathbb{X}$ as the parameter $a_{l}$ varies. As the threshold $a_{l}$ increases, more simplices are included in the filtration, revealing more complex interactions among the vertices based on their distances. This process helps to understand the shape of the data represented by the graph $G$.
@@
A persistence barcode $\textbf{bcd}_{k}(d)$, for each homological dimension $k$, is the visual representations of homological features as a multiset. They can be represented as follows:
$\textbf{bcd}_{k}(d)=\{[a,b]\mid a,b\in\mathbb{R},a<b\}.$
@@
In [30], an algorithm for computing persistence barcodes was presented. Let $\mathbf{D}$ be the boundary matrix. Each row and column of the matrix $\mathbf{D}$ correspond to the simplex, and the matrix $(\mathbf{D}_{i,j})$ is defined as $\mathbf{D}_{i,j}=\begin{cases}1&\text{ if }\sigma_{i}\in\partial\sigma_{j}\\ 0&\text{ otherwise}\end{cases}$. The algorithm performs the column operations known as reduction to obtain the reduced binary matrix $\mathbf{R}$, starting from $\mathbf{D}$. Let $\text{low}_{\mathbf{R}}(j)$ represent the row index of the last nonzero element in the $j$th column of $\mathbf{R}$, or be undefined if the $j$th column consists solely of zeros. The column operation algorithm proceeds as follows:
@@
[…]
@@
    If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $\sigma_{j}$ causes the cycle $\tau$ to die, where the death time is the value of $\sigma_{j}$. If there is no such $j$, the cycle $\tau$ will persist infinitely.
@@
    Suppose that the $i$ th column of $\mathbf{R}$ corresponding to the $k$-simplex $\sigma_{i}$ is zero. Let $f$ be the real-valued function induced by the distance $d$ as described in Section 2.2. If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ with $f(\sigma_{i})<f(\sigma_{j})$, then $[f(\sigma_{i}),f(\sigma_{j})]$ is in $\textbf{bcd}_{k}(d)$. Also, if there is no $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $[f(\sigma_{i}),\infty]$ is an element of $\textbf{bcd}_{k}(d)$.
@@
$\bigcup\{[f(\sigma_{i}),\infty]\mid\sigma_{i}\text{ is a $k$-simplex, there isno $\sigma_{j}$ such that }i=\text{low}_{\mathbf{R}}(j)\}.$
@@
Consider the ordering function $\pi$ on $\mathbb{X}$ that determines the order of simplices with the same value for each $\mathbb{X}_{a_{i}}$. We call this function as simplex ordering on $\mathbb{X}$ from the function $f$. Let $\mathbf{D}_{\pi}$ be the boundary matrix with $\pi$. Then $\mathbf{D}_{\pi}$ is uniquely determined if we fix a simplex ordering $\pi$ on $\mathbb{X}$. Also, if $\mathbf{D}_{\pi}$ is uniquely determined, then every $[f(\sigma_{i}),f(\sigma_{j})]$ in $\textbf{bcd}_{k}(d)$ is uniquely determined by $\sigma_{i}$. We obtain the set version of the persistence barcode as
@@
where $\textbf{bcd}^{\pi}_{k}(d)$ denotes the barcode with the ordering function $\pi$. For each element $[f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$ of $\textbf{bcd}^{\pi}_{k}(d)$, $\sigma_{i}$ is called the birth $k$-simplex of $[f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$.
@@ 2.3 Interrelations between two persistence barcodes
Let $G=(V,E,W_{E})$ be a connected weighted graph. Then there exists an injective function $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ defined as $\varphi([a,b])=[a,c]$ such that $b\leq c$ for any $[a,b]\in\textbf{bcd}_{1}(d_{\text{weight}})$.
@@
As illustrated in Figure 1, the presence of an inequality between the two arbitrary distances does not guarantee an injection between their $1$-dimensional persistence barcodes in general. However, Theorem 2.4 shows the existence of a pair of distance definitions that could provide an injective relationship in the $1$-dimensional barcode. For example, a possible pair would be $d_{\text{edge}}$ and $d_{\text{weight}}.$
@@
Figure 2: An example where $\textbf{bcd}_{1}(d_{\text{edge}})$ and $\textbf{bcd}_{1}(d_{\text{weight}})$ differ yet they form a hierarchical structure within $1$-dimensional hole structures. Bold lines represent newly formed edges, while dashed lines indicate edge that appear at different values.
Figure 2 shows the order in which edges appear in a Rips filtration, based on $d_{\text{edge}}$ and $d_{\text{weight}}$. When calculating $d_{\text{edge}}$, the direct path from $a$ to $b$ yields a value of $5$, but for $d_{\text{weight}}$, the path from $a$ to $b$ via $c$ results in a value of $4$. Due to this difference, $\textbf{bcd}_{1}(d_{\text{edge}})=\{[3,4],[4,5]\}$ and $\textbf{bcd}_{1}(d_{\text{weight}})=\{[3,4]\}$ differ and the injection $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ from Theorem 2.4 satisfies $\varphi([3,4])=[3,4]$.
@@ 3.1 Path-representable distance
Definition 3.2.
(Dominated distance) Consider a connected weighted graph $G=(V,E,W_{E})$ and its path-representable distance $d$ represented by the path choice function $g$ on $G$. We refer to $d$ as weight-dominated if $g$ is weight-dominated and as cost-dominated if $g$ is cost-dominated.
@@
(i) First, we show that $d_{\text{weight}}$ and $d_{\text{edge}}$ are path-representable distances. Consider $d_{\text{weight}}$. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$. Denote $p_{1,k}$ by the minimal path from $v_{1}$ to $v_{k}$ such that $d_{\text{weight}}(v_{1},v_{k})=\sum_{e\in E(p_{1,k})}W_{E}(e)$ for $k=2,3,\ldots,n$. Consider $p_{1,2}$ and $p_{1,3}$. Suppose $p_{1,2}$ and $p_{1,3}$ meet at the midpoint $z\neq v_{1}$. Then, denote the subpath from $v_{1}$ along $p_{1,2}$ to $z$ by $\alpha$, and the subpath from $v_{1}$ along $p_{1,3}$ to $z$ by $\beta$. By the minimality of $p_{1,2}$ and $p_{1,3}$, $\sum_{e\in E(\alpha)}W_{E}(e)=\sum_{e\in E(\beta)}W_{E}(e)$. This is because if one side were smaller, we could find a shorter path by replacing the other. Now, redefine $p_{1,3}$ as a new path starting at $v_{1}$, following $\alpha$ to $z$, and then from $z$ to $v_{3}$ along the path of $p_{1,3}$. Repeat this process until the midpoint where $p_{1,2}$ and $p_{1,3}$ meet disappears.
This process ensures one of the followings: (a) $p_{1,2}$ and $p_{1,3}$ meet only at $v_{1}$ or not at all, or (b) they start from $v_{1}$, overlap only on a common subpath, and do not meet anywhere else until the end. We call the above replacement process a ()(*)( ∗ )-replacement.
Next, perform the ()(*)( ∗ )-replacement for $p_{1,4}$ with respect to $\{p_{1,2},p_{1,3}\}$, and repeat this until $p_{1,n}$ is ()(*)( ∗ )-replaced with respect to $\{p_{1,2},p_{1,3},\ldots,p_{1,n-1}\}$. Then, take $p_{2,1}$, reverse its direction to be the same as $p_{1,2}$, and ()(*)( ∗ )-replace $p_{2,3}$ with respect to $p_{2,1}$. Continue to $p_{2,4}$ with respect to $\{p_{2,1},p_{2,3}\}$, and repeat this process until $p_{2,n}$ is ()(*)( ∗ )-replaced with respect to $\{p_{2,1},p_{2,3},\ldots,p_{2,n-1}\}$. By continuing this process, we ensure that for any distinct $i$ and $j$, the consistency property for $p_{i,j}$ is satisfied. Finally, define the path choice function $g$ as $g(v_{i},v_{j})=p_{i,j}$, thus proving that $d_{\text{weight}}(v_{i},v_{j})=\sum_{e\in g(v_{i},v_{j})}W_{E}(e)$
The ()(*)( ∗ )-replacement process can also be applied in the case of $d_{\text{edge}}$. If there is a midpoint $z$, both the number of edges in the two different paths with the same starting points and the sum of the weights up to the midpoint must be the same. Thus the replacement can be done similarly. Therefore, $d_{\text{edge}}$ becomes the path-representable distance by undergoing the similar procedure.

### v2

# abEunwoo Heo these authors contributed equally to this work    abByeongchan Choi22footnotemark: 2    abJae-Hun Jung Corresponding author

@@ Abstract
Persistent homology (PH) has been widely applied to graph data to extract topological features. However, little attention has been paid to how different distance functions on a graph affect the resulting persistence barcodes and their interpretations. In this paper, we define a class of distances on graphs, called path-representable distances, and investigate structural relationships between their induced persistent homologies. In particular, we identify a nontrivial injection between the $1$-dimensional barcodes induced by two commonly used graph distances: the unweighted and weighted shortest-path distances. We formally establish sufficient conditions under which such embeddings arise, focusing on a subclass we call cost-dominated distances. The injection property is shown to hold in 0- and $1$-dimensions, while we provide counterexamples for higher-dimensional cases. To make these relationships measurable, we introduce the total persistence difference (TPD), a new topological measure that quantifies changes between filtrations induced by cost-dominated distances on a fixed graph. We prove a stability result for TPD when the distance functions admit a partial order and apply the method to the SNAP EU Research Institution E-Mail dataset. TPD captures both periodic patterns and global trends in the data, and shows stronger alignment with classical graph statistics compared to an existing PH-based measure applied to the same dataset.
@@ 1 Introduction
Topological data analysis (TDA), developed recently, has proven its usefulness in various applications [14, 11, 39, 17]. Persistent homology (PH) is one of the primary tools in TDA, using homological characteristics of the data at various scales [11]. The PH method involves constructing the complex in different scales sequentially, capturing how homology changes as the complex is being built and refined. This procedure is known as filtration. These hierarchical changes of homology enable the PH method to provide the topological inference of the data.
@@
Once the appropriate transform is identified, applying the PH method requires defining a proper metric for the transformed geometric object, determining the distance between points in the point cloud or between nodes in the graph representation. There is no single way to define this metric or distance as there can be multiple definitions. Definitions of such distances have been empirically established in different fields of study, yet there is a lack of research on how these varying definitions of distance impact PH calculations. For a brief functorial overview of common graph-to-persistence pipelines in TDA, see Appendix A.
@@
On the other hand, the shortest path as the path connecting two vertices with the least number of edges can also be considered. In a similar sense, there are data in which the relevancy between two nodes is more explained in the path passing through the minimum number of edges than in that with minimum sum of weights. In such a case, it is natural to define the distance between two vertices using such paths. The first approach was conducted in previous research [37]. We canonically reformulated the definition of distance $d_{\text{edge}}(v,w)$ between two points on the graph to be the minimum sum of weights between the paths connecting $v$ and $w$ with the minimum number of edges.
@@
In this research we found an interesting relation between $d_{\text{weight}}$ and $d_{\text{edge}}$. That is, for $d_{\text{weight}}\leq d_{\text{edge}}$, we have $\textbf{bcd}_{1}(d_{\text{weight}})\subseteq\textbf{bcd}_{1}(d_{\text{edge}})$, where $\textbf{bcd}_{1}$ is the $1$-dimensional barcode. This property is quite unusual and is generally not expected to hold, as depicted in Fig. 1. To explore this, we introduce a more general concept of distance, termed path-representable distance, and identify a class of cost-dominated distances for which this inclusion always holds. Furthermore, in this paper, we prove that such an injection relation holds among any path-representable cost-dominated distances for $1$-dimensional homology.
Beyond the existence of this injective structure, it is natural to ask how much the $1$-dimensional persistence changes as the distance is altered. Given two cost-dominated path-representable distances $d_{i}\leq d_{j}$ on the same graph, we define the total persistence difference (TPD) as a nonnegative quantity that compares the corresponding $1$-dimensional barcodes cycle by cycle. Each birth edge that appears for both distances contributes the increase in its death value under $d_{j}$ relative to $d_{i}$, and each birth edge that appears only for $d_{j}$ contributes the persistence of the associated interval. In this way TPD measures how the $1$-dimensional structure extends when the distance on the graph is changed.
We prove that TPD enjoys a Lipschitz stability with respect to the pointwise difference between the underlying distances. More precisely, for cost-dominated path-representable distances $d_{i}\leq d_{j}$ we obtain an upper bound of the form
$\mathrm{Diff}_{p}\bigl(\textbf{bcd}_{1}(G,d_{j}),\,\textbf{bcd}_{1}(G,d_{i})\bigr)\leq\gamma\,\|d_{j}-d_{i}\|_{\infty},$
where the constant $\gamma$ depends only on the number of vertices and edges of the graph.
Finally, we investigate how TPD behaves on real temporal graph data. We apply our framework to the SNAP EU Research Institution E-mail dataset, where for each day we build a weighted graph from all e-mails exchanged during that day and consider two natural path-representable distances on this graph. On this dataset TPD is typically close to zero on structurally simple days, and becomes larger when the daily graph has higher degree and clustering. We also compare TPD with a previously proposed PH-based measure computed from overlapping time windows by reproducing the pipeline of [24]. In this comparison, TPD shows stronger Pearson correlations with basic graph statistics and better reflects periods of low activity, while the Wasserstein-based day-to-day distance does not decay on weekends where the graph structure is sparse or trivial.
@@
This paper is composed of the following sections. In Section 2, we explain distinct distance definitions over the given graph, particularly the weight- and edge-minimal distances and the possible relations among them. In Section 3, we introduce the concept of path-representable distance, a generalization of $d_{\text{weight}}$ and $d_{\text{edge}}$ and we introduce the main theory related to path-representable distances. In Section 4, we rigorously prove the main propositions and theorem. In Section 5, we make an experiment to explain how the main theorem works in higher dimensions. In Section 6, we define the total persistence difference and establish its stability theorem. In Section 7, we apply our method to the e-mail dataset and compare TPD with existing PH-based quantities. Finally, in Section 8, we provide a brief concluding remark.
@@ 2.1 Different distance definitions
[…]
@@ 2.2 Persistent homology
for any $k$-simplex $\sigma\in\mathbb{X}$ whenever $k\geq 1$, otherwise $f(\sigma)=0$. For a sequence of real numbers $0=a_{0}\leq a_{1}\leq\ldots\leq a_{m}$, if we denote $\mathbb{X}_{a_{l}}\coloneqq f^{-1}((-\infty,a_{l}])$, then we have a filtration
@@
Note that $\mathbb{X}_{a_{0}}=V$ and $\mathbb{X}_{a_{m}}=\mathcal{P}(V)$. This filtration is called the Vietoris–Rips (Rips) filtration. The filtration induced by $f$ allows us to study topological changes in the simplicial complex $\mathbb{X}$ as the parameter $a_{l}$ varies. As the threshold $a_{l}$ increases, more simplices are included in the filtration, revealing more complex interactions among the vertices based on their distances. This process helps to understand the shape of the data represented by the graph $G$.
@@
A persistence barcode $\textbf{bcd}_{k}(d)$, for each homological dimension $k$, is a visual representation of homological features as a multiset. They can be represented as follows:
$\textbf{bcd}_{k}(d)=\{[\beta,\delta]\mid\beta,\delta\in\mathbb{R},\beta<\delta\}.$
@@
In [15], an algorithm for computing persistence barcodes was presented. Let $\mathbf{D}$ be the boundary matrix. Each row and column of the matrix $\mathbf{D}$ corresponds to a simplex, and the matrix $(\mathbf{D}_{i,j})$ is defined as $\mathbf{D}_{i,j}=\begin{cases}1&\text{ if }\sigma_{i}\in\partial\sigma_{j}\\ 0&\text{ otherwise}\end{cases}$. The algorithm performs the column operations known as reduction to obtain the reduced binary matrix $\mathbf{R}$, starting from $\mathbf{D}$. Let $\text{low}_{\mathbf{R}}(j)$ represent the row index of the last nonzero element in the $j$th column of $\mathbf{R}$, or be undefined if the $j$th column consists solely of zeros. The column operation algorithm proceeds as follows:
@@
1: Define $\mathbf{R}\coloneqq\mathbf{D}\coloneqq\text{the boundary matrix}$
@@
    If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $\sigma_{j}$ causes the cycle $\tau$ to die, where the death value is the value of $\sigma_{j}$. If there is no such $j$, the cycle $\tau$ will persist infinitely.
@@
    Suppose that the $i$ th column of $\mathbf{R}$ corresponding to the $k$-simplex $\sigma_{i}$ is zero. Let $f$ be the real-valued function induced by the distance $d$ as described in Section 2.2. If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ with $f(\sigma_{i})<f(\sigma_{j})$, then $f(\sigma_{i}),f(\sigma_{j})]$ is in $\textbf{bcd}_{k}(d)$. Also, if there is no $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $f(\sigma_{i}),\infty]$ is an element of $\textbf{bcd}_{k}(d)$.
@@
[…]
@@
Consider the ordering function $\pi$ on $\mathbb{X}$ that determines the order of simplices with the same value for each $\mathbb{X}_{a_{i}}$. We call this function the simplex ordering on $\mathbb{X}$ from the function $f$. Let $\mathbf{D}_{\pi}$ be the boundary matrix with $\pi$. Then $\mathbf{D}_{\pi}$ is uniquely determined if we fix a simplex ordering $\pi$ on $\mathbb{X}$. Also, if $\mathbf{D}_{\pi}$ is uniquely determined, then every $f(\sigma_{i}),f(\sigma_{j})]$ in $\textbf{bcd}_{k}(d)$ is uniquely determined by $\sigma_{i}$. We obtain the set version of the persistence barcode as
@@
where $\textbf{bcd}^{\pi}_{k}(d)$ denotes the barcode with the ordering function $\pi$. For each element $f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$ of $\textbf{bcd}^{\pi}_{k}(d)$, $\sigma_{i}$ is called the birth $k$-simplex of $f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$.
@@ 2.3 Interrelations between two persistence barcodes
Let $G=(V,E,W_{E})$ be a connected weighted graph. Then there exists an injective function $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ defined as $\varphi([\beta,\delta])=[\beta,\delta^{\prime}]$ such that $\delta\leq\delta^{\prime}$ for any $\beta,\delta]\in\textbf{bcd}_{1}(d_{\text{weight}})$.
@@
As illustrated in Fig. 1, the presence of an inequality between the two arbitrary distances does not guarantee an injection between their $1$-dimensional persistence barcodes in general. However, Theorem 2.4 shows the existence of a pair of distance definitions that could provide an injective relationship in the $1$-dimensional barcode. For example, a possible pair would be $d_{\text{edge}}$ and $d_{\text{weight}}.$
@@
Figure 2: An example where $\textbf{bcd}_{1}(d_{\text{edge}})$ and $\textbf{bcd}_{1}(d_{\text{weight}})$ differ yet they form a hierarchical structure within 11-dimensional hole structures. Bold lines represent newly formed edges, while dashed lines indicate edges that appear at different values.
Fig. 2 shows the order in which edges appear in a Rips filtration, based on $d_{\text{edge}}$ and $d_{\text{weight}}$. When calculating $d_{\text{edge}}$, the direct path from $a$ to $b$ yields a value of $5$, but for $d_{\text{weight}}$, the path from $a$ to $b$ via $c$ results in a value of $4$. Due to this difference, $\textbf{bcd}_{1}(d_{\text{edge}})=\{[3,4],[4,5]\}$ and $\textbf{bcd}_{1}(d_{\text{weight}})=\{[3,4]\}$ differ and the injection $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ from Theorem 2.4 satisfies $\varphi([3,4])=[3,4]$.
@@ 3.1 Path-representable distance
Definition 3.2 (Dominated distance).
Consider a connected weighted graph $G=(V,E,W_{E})$ and its path-representable distance $d$ represented by the path choice function $g$ on $G$. We refer to $d$ as weight-dominated if $g$ is weight-dominated and as cost-dominated if $g$ is cost-dominated.
@@
(i) First, we show that $d_{\text{weight}}$ and $d_{\text{edge}}$ are path-representable distances. Consider $d_{\text{weight}}$. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$. Denote $p_{1,k}$ by the minimal path from $v_{1}$ to $v_{k}$ such that $d_{\text{weight}}(v_{1},v_{k})=\sum_{e\in E(p_{1,k})}W_{E}(e)$ for $k=2,3,\ldots,n$. Consider $p_{1,2}$ and $p_{1,3}$. Suppose $p_{1,2}$ and $p_{1,3}$ meet at the midpoint $z\neq v_{1}$. Then, denote the subpath from $v_{1}$ along $p_{1,2}$ to $z$ by $p_{1,2}|_{z}$, and the subpath from $v_{1}$ along $p_{1,3}$ to $z$ by $p_{1,3}|_{z}$. By the minimality of $p_{1,2}$ and $p_{1,3}$, $\sum_{e\in E(p_{1,2}|_{z})}W_{E}(e)=\sum_{e\in E(p_{1,3}|_{z})}W_{E}(e)$. This is because if one side were smaller, we could find a shorter path by replacing the other. Now, redefine $p_{1,3}$ as a new path starting at $v_{1}$, following $p_{1,2}|_{z}$ to $z$, and then from $z$ to $v_{3}$ along the path of $p_{1,3}$. Repeat this process until the midpoint where $p_{1,2}$ and $p_{1,3}$ meet disappears.
This process ensures one of the following: (a) $p_{1,2}$ and $p_{1,3}$ meet only at $v_{1}$ or not at all, or (b) they start from $v_{1}$, overlap only on a common subpath, and do not meet anywhere else until the end. We call the above replacement process a ()(*)-replacement.
Next, perform the ()(*)-replacement for $p_{1,4}$ with respect to $\{p_{1,2},p_{1,3}\}$, and repeat this until $p_{1,n}$ is ()(*)-replaced with respect to $\{p_{1,2},p_{1,3},\ldots,p_{1,n-1}\}$. Then, take $p_{2,1}$, reverse its direction to be the same as $p_{1,2}$, and ()(*)-replace $p_{2,3}$ with respect to $p_{2,1}$. Continue to $p_{2,4}$ with respect to $\{p_{2,1},p_{2,3}\}$, and repeat this process until $p_{2,n}$ is ()(*)-replaced with respect to $\{p_{2,1},p_{2,3},\ldots,p_{2,n-1}\}$. By continuing this process, we ensure that for any distinct $i$ and $j$, the consistency property for $p_{i,j}$ is satisfied. Finally, define the path choice function $g$ as $g(v_{i},v_{j})=p_{i,j}$, thus proving that $d_{\text{weight}}(v_{i},v_{j})=\sum_{e\in g(v_{i},v_{j})}W_{E}(e)$ .
The ()(*)-replacement process can also be applied in the case of $d_{\text{edge}}$. If there is a midpoint $z$, both the number of edges in the two different paths with the same starting points and the sum of the weights up to the midpoint must be the same. Thus the replacement can be done similarly. Therefore, $d_{\text{edge}}$ becomes the path-representable distance by undergoing the similar procedure.